# Does AI Code Review Lead to Code Changes? A Case Study of GitHub Actions

### v1

@@ IV-C RQ3: Which factors impact the likelihood that code review comments lead to code changes?
Approach: To answer RQ3, we conducted an interpretable analysis of factors influencing whether code review comments lead to code changes. We first applied the optimal approach established in RQ2 to classify all 5,652 review comments in our refined addressing dataset, determining their validity and addressing status. To overcome the inherent lack of interpretability in LLMs, we engineered a structured feature set covering various dimensions that potentially affect response behavior. We then trained a Random Forest classifier to fit the LLM-derived addressing results and used SHAP (SHapley Additive exPlanations) [22] analysis to interpret the influence and directionality of each feature on the model’s predictions.
TABLE IX: RQ3: 45 Considered Features Across Four Dimensions (36 Retained After AutoSpearman’s Redundant Feature Removal).
@@
Source Features (10): Who generated the comment and how it was generated Is_Human 1 if written by a human; 0 if AI-generated Is_File_Level_Action 1 if generated by file-level action; 0 otherwise Action_[4 actions] One-hot (4): Which action produced the comment Trigger_[auto/manual] One-hot (2): Action trigger type: Auto/Manual (all-zero if humans) LLM_[GPT-3.5/4] One-hot (2): LLM used for reviewing (all-zero if unknown/humans) Repository Features (5): Repository-level context where the review occurred Repo_File_Count Total number of files in the repository Repo_File_Size Total size of files in the repository (in bytes) Repo_Issue_Count Total number of issues in the repository Repo_PR_Count Total number of PRs in the repository Repo_Contributor_Count Number of non-bot contributors in the repository Modification Features (17): Characteristics of the code changes under reviewed Author_Is_Bot 1 if the author of reviewed commit is a bot; 0 otherwise Author_Is_Anon 1 if the author is anonymous; 0 otherwise Author_Prior_Commits Number of previous commits by the author to this repository Commit_Changed_File Number of files changed in the commit (vs. PR base commit) Commit_Change/Add/Del Total lines changed/added/deleted in changed files Commit_Base_Lines Total lines in all changed files at base commit File_Is_Code 1 if a certain reviewed file is a programming file (via Linguist) File_Depth Directory nesting level of the file File_Add/Del/Change Lines added/deleted/changed in the reviewed file File_Base_Lines Line count of the file at base commit Comment_Add/Del/Change Lines added/deleted/changed the comment scope Comment Features (13): Structure and content of the review comment itself Timeline_Index Position of comment in chronological order within PR Prior_Comment_Len Total character length of all prior comments in the PR Inline_Code 1 if contains inline code snippets (`code`) Multiline_Code 1 if contains multiline code blocks (```code```) Text_Length Total Length of the current comment (in characters) Code_Length Total length of all code snippets among the comment Code_Text_Ratio Ratio of code length to total comment length LDA_Topic_[6] Six LDA topic probabilities with GPT-4.1 labeled themes: 0: Dependency Version Update and Compatibility; 1: UI Usability and Accessibility Enhancements; 2: Concurrency Control and Locking Robustness; 3: Module Import Optimization; 4: Variable Naming and Code Readability; 5: Error Handling and Code Documentation;
Phase I: LLM-Based Labeling of Comment Addressing. Following the best-performing setup on the annotated dataset, we first used gpt-4.1 to assess the validity of all 5,652 comments, identifying 3,955 as valid. Subsequently, we used openai-o3-mini to evaluate their addressing status. Given that 65.3% of comments in our dataset had no subsequent modifications (and we were therefore able to trivially classify them as “Not Addressed” for Stage-2), our approach can achieve approximately 95% accuracy in identifying valid and addressed comments across the entire dataset. The complete labeling results are shown in Table VIII. There is a discrepancy in annotation distributions between Table VI (manual labels) and Table VIII (automated labels). We further explain this in the results of RQ3. For the subsequent analysis, we focus on the 3,879 valid comments with a definitive addressing label (further excluding 76 uncertain cases) to examine what factors influence whether they finally lead to the code changes. We discuss the potential causes of invalid comments and possible mitigation strategies in Section V.
Phase II: Structured Feature Engineering. Inspired by previous works [9, 23], We initially built a 45-feature set across four dimensions to comprehensively capture the possible impacting factors (Table IX). Specifically, Source Features capture who generated the comment (a human developer or one of four actions) and two key aspects of action configuration: the trigger mode (automatic vs. manual) and the LLM used (GPT-3.5 vs. GPT-4). These configurations were extracted by parsing the workflow file at the comment’s associated commit. If the workflow included an ‘if’ conditional trigger, we labeled the comment as manually triggered; otherwise, it was considered automatic. For LLM selection (all four actions in our final analysis employed OpenAI models), we mapped explicitly specified models to either the GPT-3.5 or GPT-4 family; if unspecified, the model was marked as “unknown”. Repository Features reflect the project context, including overall size and activity level (e.g., number of issues). Modification Features characterize the code changes under review, including the authorship and change properties. We used Author_Prior_Commits to estimate the author’s familiarity with the project. Edge cases, such as when the author was identified as a bot or an anonymous user (1.8% of valid comments), were flagged using Author_Is_Bot/Anon, with Author_Prior_Commits set to 0 accordingly. We further described change scope at the commit, file, and comment levels, along with the volume of the involved files before changes were made. Comment Features reflect the characteristics of the comment itself, including structural features (e.g., its position in the PR discussion thread) and textual features (e.g., code ratio). To capture comment intent, we applied Latent Dirichlet Allocation (LDA) [24] to the 3,879 comments. Following hyperparameter tuning strategies in prior work [25], we selected six as the optimal number of topics. We then used gpt-4.1 to summarize the top ten comments for each topic to generate interpretable labels (as shown in Table IX). To further address potential multicollinearity, we applied AutoSpearman correlation analysis [26] ($\rho\!>\!0.7$ threshold) and removed 9 redundant features (marked with strikethrough), yielding a final set of 36 features.
TABLE X: RQ3: Top features for comment addressing, grouped and ranked by group-wise SHAP Importance ($\sum|\phi|$)).
@@
Feature Importance ($|\phi|$) Directionality ($\rho$) Source Feature (10): $\sum|\phi|$=0.1646; $\mu|\phi|$=0.0165 Is_Human 0.0549 0.99 Trigger_auto 0.0459 -0.97 Is_File_Level_Action 0.0400 -0.96 Comment Feature (10): $\sum|\phi|$=0.1059; $\mu|\phi|$=0.0096 Code_Text_Ratio 0.0159 0.89 Text_Length 0.0148 -0.28 LDA_Topic_1 0.0119 0.42 LDA_Topic_2 0.0116 0.55 Modification Feature (12): $\sum|\phi|$=0.0756; $\mu|\phi|$=0.0063 Author_Prior_Commits 0.0243 -0.67 Repository Feature (3): $\sum|\phi|$=0.0422; $\mu|\phi|$=0.0141 Repo_File_Size 0.0228 0.73 Repo_Issue_Count 0.0114 0.67
@@
Phase III: Model Fitting and Interpretation. Since our feature set spans four dimensions, we chose a Random Forest classifier with SHAP explanations to better model and interpret the potential complex, non-linear interactions between features and whether a comment is addressed. We further discuss this choice in Section VI. To enable focused analysis, we mapped the 3,879 comments into a binary classification task: comments labeled as “Valid-Partially” or “Valid-Fully” were grouped as the “Addressed” class (25.1%), with “Valid-Unaddressed” comments forming the “Not Addressed” class (74.9%). Upon our refined 36 features, we trained a Random Forest classifier (80% training data) that achieved 88.5% overall accuracy (Macro-F1=0.854) on the test set, indicating that our engineered features are effective predictors of comment addressing. Since Random Forest’s native feature importance metrics lack directional insights, we applied SHAP to quantify the feature’s contribution towards “Addressed” predictions. SHAP is a widely-used technique to interpret learning-based models. It explains individual predictions by assigning each feature a Shapley value, which represents its average marginal contribution to the prediction across all possible feature combinations. For each feature, we calculated two interpretability metrics: 1) SHAP Importance: the mean absolute SHAP value across test instances, reflecting global predictive importance; and 2) SHAP Directionality: the Pearson correlation between feature values and their SHAP values, indicating the direction of association (positive or negative) with “Addressed” predictions.
Results: The distribution shift between Table VI and Table VIII is reasonable and stems from our sampling strategy. For the manual annotation, we purposefully sampled only those comments where the reviewed file was subsequently modified. This was a necessary precondition for assessing if a comment was addressed, but it naturally biased the sample toward “Valid” and “Addressed” classifications. In contrast, the automated annotation was applied to the entire comment dataset, which included a large volume of comments (65.3%) with no subsequent file changes. Following our procedure, if these comments were identified as “Valid” in Stage-1, they were automatically classified as “Not Addressed” in Stage-2. Consequently, the proportion of “Valid, Addressed” comments is lower in the automated setting.
Table X presents the top 10 features with SHAP Importance exceeding the average (0.0108 across 36 features), organized by feature groups. For each group, we also report the total SHAP Importance ($\sum|\phi|$) and average SHAP Importance ($\mu|\phi|$), and rank them by $\sum|\phi|$. The prominent ranking of “Source features” and “Comment features” suggests that the configuration and formatting of automated reviews are strongly associated with whether comments are addressed.
@@
Notably, Source Features represent the most significant group, showing a clear difference in developers’ responses based on the comment origins (human vs. actions; file-level vs. hunk-level). The negative correlation for Is_File_Level_Action ($\rho\!=\!-0.96$) aligns with the trend in Table VIII, where hunk-level review actions generally outperform file-level counterparts. Specifically, coderabbitai/ai-pr-reviewer achieved the highest valid and addressed rate (19.2%). Automatic triggers also show a negative correlation with addressing (Trigger_auto, $\rho\!=\!-0.97$). To detail this effect, we analyzed data distribution by trigger modes across the four actions (Table VIII). We found that in the two actions with manually triggered comments (Action ID-1 and ID-2), manually triggered comments consistently showed higher addressing rates. Interestingly, the LLM choice shows a low predictive importance (LLM_GPT-4 ranked 31/36, LLM_GPT-3.5 ranked 28/36 in the feature list). This might be due to the “unknown” LLM selection for some comments, potentially diluting its perceived importance. However, when an LLM was explicitly identified (Table XI), GPT-4 generated comments tended to have better addressing rates.
TABLE XI: RQ3: Addressing rates by trigger mode and LLM series across review actions (p-values from Fisher’s Exact Test)
Action Trigger Type Total Addressed (%) p-value LLM Series Total Addressed (%) p-value Total 4 Auto 2,886 9.0% p>0.05p\!>\!0.05 GPT-4 922 8.4% $\boldsymbol{p\!\leq\!0.05}$ Manual 104 14.4% GPT-3.5 1,025 4.4% ID-1 Auto 1,595 6.8% $\boldsymbol{p\!\leq\!0.05}$ GPT-4 236 16.9% $\boldsymbol{p\!\leq\!0.05}$ Manual 86 12.8% GPT-3.5 1,021 4.4% ID-2 Auto 602 0.5% $\boldsymbol{p\!\leq\!0.05}$ GPT-4 621 1.1% Manual 18 22.2% ID-3 Auto 534 25.7% GPT-4 48 60.4% $\boldsymbol{p\!\leq\!0.05}$ Manual GPT-3.5 4 0.00% ID-4 Auto 155 7.1% GPT-4 18 5.6% Manual
TABLE XII: RQ3: Comment Addressing rates by binned Code_Text_Ratio and Author_Prior_Commits (Human vs. All Four Actions). Each cell shows the comment count in the bin and its addressed ratio.
Bins of Code_Text_Ratio Human Four Actions Bins of Author_Prior_Commits Human Four Actions Total (Addressed%) Total (Addressed%) (–0.00, 0.00] 320 (71.2%) 642 (4.2%) (–0.00, 30] 558 (80.1%) 218 (16.1%) (0.00, 0.11] 61 (65.6%) 669 (7.2%) (30, 124] 140 (82.1%) 639 (16.1%) (0.11, 0.26] 95 (75.8%) 635 (5.7%) (124, 319] 131 (71.8%) 642 (12.1%) (0.26, 0.52] 107 (78.5%) 621 (10.5%) (319, 1,013] 32 (81.2%) 763 (4.5%) (0.52, 1.00] 306 (89.9%) 423 (23.2%) (1013, 4,316] 28 (60.7%) 728 (3.3%)
@@
For Comment features, we observe that the comments were more likely to be addressed if they were concise (Text_Length, $\rho\!=\!-0.28$) or contained a higher proportion of code (Code_Text_Ratio, $\rho\!=\!0.89$), especially when they included multi-line code blocks(Has_Multiline_Code, $\rho\!=\!0.78$ not shown in the table). As shown in Table XII (left), the addressing rate for both human and AI-generated comments increased noticeably when Code_Text_Ratio exceeded 0.5. This may be because multi-line code is often used to illustrate concrete solutions, allowing developers to follow comments with minimal effort through copy-pasting. Regarding the comment topic, comments were addressed more often when they pertained to “UI Usability and Accessibility Enhancements” (e.g., adding aria-label for screen reader support) (LDA_Topic_1, $\rho\!=\!0.42$), as well as “Concurrency Control and Locking Robustness” (e.g., using a ‘set-if-not-exists’ pattern) (LDA_Topic_2, $\rho\!=\!0.55$).
@@
Within Modification Features, review comments targeting commits by experienced contributors were less likely to be addressed (Author_Prior_Commits, $\rho\!=\!-0.67$), suggesting that AI-generated reviews may be more helpful for newcomers. As shown in Table XII (right), valid comments directed at project newcomers ($\textit{Author\_Prior\_Commits}\!\leq\!124$) achieved 16% addressing rate, compared to just 3.3% for the most experienced contributors. Notably, human reviews mainly focused on these newcomers, with 79% of comments directed at them, aligning with typical community practices. In contrast, automated actions reviewed code changes regardless of author experience, which may help mitigate potential experience blindspots in the review process. Finally, positive correlations for Repository Features demonstrated that comments were more likely to be addressed in large and active projects (Repo_File_Size, $\rho\!=\!0.73$; Repo_Issue_Count, $\rho\!=\!0.67$).
@@
Answer to RQ3: Comment addressing is strongly associated with Source Features (human vs. AI and specific action type), followed by Comment Features (conciseness is better), highlighting the importance of thoughtful automated review design.

### v2

@@ IV-C RQ3: Which factors impact the likelihood that code review comments lead to code changes?
Approach: To answer RQ3, we conducted an interpretable analysis to identify the factors influencing whether review comments lead to code changes. We focus specifically on the 3,879 valid comments from RQ2 that have a definitive addressing label (excluding 76 uncertain cases), examining the factors that determine whether they are addressed. For invalid comments, we discuss their potential causes and mitigation strategies in Section V. To quantify the impact of different factors, we engineered a structured feature set covering various dimensions that potentially affect response behavior. Since direct interpretation of LLM decision-making is challenging, we trained a Random Forest classifier to fit the addressing results derived in RQ2 and employed SHAP (SHapley Additive exPlanations) [34] analysis to interpret the influence and directionality of each feature on the model’s predictions.
TABLE X: RQ3: 36 Selected Features Across Four Dimensions Used for Analyzing Factors Impacting Comment Addressing.
@@
Source Features (10): Who generated the comment and how it was generated Is_Human 1 if written by a human; 0 if AI-generated Is_File_Level_Action 1 if generated by file-level action; 0 otherwise Is_Action_[4] One-hot (4): Which action produced the comment Trigger_[auto/manual] One-hot (2): Action trigger type: Auto/Manual (all-zero if human) LLM_[GPT-3.5/4] One-hot (2): LLM used for reviewing (all-zero if human) Repository Features (3): Repository-level context where the review occurred Repo_File_Size Total size of files in the repository (in bytes) Repo_Issue_Count Total number of issues in the repository Repo_Contributor_Count Number of non-bot contributors in the repository Modification Features (12): Characteristics of the code changes under reviewed Author_Is_Bot 1 if the author of reviewed commit is a bot; 0 otherwise Author_Is_Anon 1 if the author is anonymous; 0 otherwise Author_Prior_Commits Number of previous commits by the author to this repository Commit_Del Total lines deleted in changed files Commit_Base_Lines Total lines in all changed files at base commit File_Is_Code 1 if a certain reviewed file is a programming file (via Linguist) File_Depth Directory nesting level of the file File_Add/Del Lines added/deleted in the reviewed file File_Base_Lines Line count of the file at base commit Comment_Add/Del Lines added/deleted the comment scope Comment Features (11): Structure and content of the review comment itself Prior_Comment_Len Total character length of all prior comments in the PR Inline_Code 1 if contains inline code snippets (`code`) Multiline_Code 1 if contains multiline code blocks (```code```) Text_Length Total Length of the current comment (in characters) Code_Text_Ratio Ratio of code length to total comment length LDA_Topic_[6] Six LDA topic probabilities with GPT-4.1 labeled themes: 0: Dependency Version Update and Compatibility; 1: UI Usability and Accessibility Enhancements; 2: Concurrency Control and Locking Robustness; 3: Module Import Optimization; 4: Variable Naming and Code Readability; 5: Error Handling and Code Documentation;
[…]
TABLE XI: RQ3: top 10 features influencing total comment addressing. Gray-shaded rows denote features that also rank among the top 10 for AI comment addressing (Table XII).
@@
Feature Rank Importance ($|\phi|$) Directionality ($\rho$) Source Feature (10): $\sum|\phi|$=0.1589; $\mu|\phi|$=0.0159 Is_Human 1 0.0449 0.99 Trigger_auto 2 0.0410 -0.96 Is_File_Level_Action 3 0.0346 -0.95 Repository Feature (3): $\sum|\phi|$=0.0342; $\mu|\phi|$=0.0114 Repo_File_Size 5 0.0184 0.70 Modification Feature (12): $\sum|\phi|$=0.0801; $\mu|\phi|$=0.0067 Author_Prior_Commits 4 0.0235 -0.69 Commit_Del 9 0.0105 -0.15 Comment Feature (11): $\sum|\phi|$=0.0959; $\mu|\phi|$=0.0010 Code_Text_Ratio 6 0.0155 0.89 Text_Length 7 0.0133 -0.24 LDA_Topic_2 8 0.0116 0.50 LDA_Topic_1 10 0.0102 0.41
@@
Phase II: Model Fitting and Interpretation. Since our feature set spans four dimensions, we chose a Random Forest classifier with SHAP explanations to better model and interpret the complex, non-linear interactions between features and comment addressing. We further discuss this choice in Section VI. To enable focused analysis, we mapped the 3,879 comments into a binary classification task: comments labeled as “Valid-Partially” or “Valid-Fully” were grouped as the “Addressed” class (25.1%), with “Valid-Unaddressed” comments forming the “Not Addressed” class (74.9%). Upon our refined 36 features, we trained a Random Forest classifier using stratified 5-fold cross-validation. This approach ensures that our performance metrics and interpretability results are robust and not biased by a single random data split. The model achieved an average of 88.6% overall accuracy (Macro-F1=0.846) across the five test folds, indicating that our engineered features are effective predictors of comment addressing. Since Random Forest’s native feature importance metrics lack directional insights, we applied SHAP to quantify each feature’s contribution towards “Addressed” predictions. Specifically, we computed SHAP values for every instance in the test set of each fold and aggregated them to form a comprehensive explanation for the entire dataset. SHAP is a widely-used technique to interpret learning-based models. It explains individual predictions by assigning each feature a Shapley value, which represents its average marginal contribution to the prediction across all possible feature combinations. Based on the aggregated SHAP values, we calculated two interpretability metrics for each feature: 1) SHAP Importance: the mean absolute SHAP value, reflecting global predictive importance; and 2) SHAP Directionality: the Pearson correlation between feature values and their SHAP values, indicating the association direction (positive or negative) with “Addressed” predictions.
Phase III: AI-Specific Modeling and Interpretation. In addition to modeling the entire dataset, we also conducted an AI-specific modeling and interpretation to ensure that the specific factors influencing the effectiveness of AI-generated comments were not overshadowed by the human-authored ones. To achieve this, we re-trained the Random Forest classifier using only the 2,990 AI-generated comments, removed the Is_Human feature, and mapped two sets of one-hot features (Trigger_[auto/manual] and LLM_[GPT-3.5/4]) into two binary features (Is_Trigger_Manual and Is_Model_GPT4). The resulting AI-Specific Model achieved an overall accuracy of 91.7% and a Macro-F1 score of 0.647 in 5-fold cross-validation. Given the highly imbalanced nature of the dataset (only 9.2% addressed comments), we consider this performance reasonable for supporting the subsequent interpretation analysis. Finally, we applied the same SHAP-based analysis to the remaining 33 features to interpret the model’s predictions.
TABLE XII: RQ3: Top 10 Features Influencing AI Comment Addressing. Gray-shaded rows denote features that also rank among the top 10 for total comment addressing (Table XI).
Feature Rank Importance ($|\phi|$) Directionality ($\rho$) Source Features (7): $\sum|\phi|$=0.0418; $\mu|\phi|$=0.0060 Is_Action_3 2 0.0177 0.92 Is_File_Level_Action 7 0.0096 -0.88 Repository Features (3): $\sum|\phi|$=0.0203; $\mu|\phi|$=0.0068 Repo_File_Size 6 0.0101 0.42 Modification Features (12): $\sum|\phi|$=0.0767; $\mu|\phi|$=0.0064 Author_Prior_Commits 1 0.0212 -0.67 Commit_Base_Lines 4 0.0111 0.11 Commit_Del 5 0.0107 -0.17 Comment_Add 8 0.0095 0.81 Comment Features (11): $\sum|\phi|$=0.0709; $\mu|\phi|$=0.0064 Code_Text_Ratio 3 0.0124 0.78 LDA_Topic_5 9 0.0085 0.61 Prior_Comment_Len 10 0.0075 0.65
Results: Table XI presents the interpretability results based on the complete dataset (including both human and AI comments). We list the top 10 most important features, organized by their feature groups. For each group, we also report the total SHAP Importance ($\sum|\phi|$) and average SHAP Importance ($\mu|\phi|$). In our subsequent analysis, we mainly focus on features that show a meaningful correlation (i.e. $|\rho|\!>\!0.3$). Notably, “Source Features” and “Comment Features” rank first and second in total importance ($\sum|\phi|$), suggesting that both who provides the feedback (source credibility) and what the feedback contains (content quality) are strongly associated with whether comments are addressed.
@@
Source Features represent the most significant group, showing a clear difference in developers’ responses based on the comment origins (human vs. actions; file-level vs. hunk-level). The negative correlation for Is_File_Level_Action ($\rho\!=\!-0.95$) aligns with the trend in Table IX, where hunk-level review actions generally outperform file-level counterparts. Specifically, coderabbitai/ai-pr-reviewer achieved the highest valid and addressed rate (19.2%). Automatic triggers also show a negative correlation with addressing (Trigger_auto, $\rho\!=\!-0.96$). To detail this effect, we analyzed data distribution by trigger modes across the four actions (Table XVI). We found that in the two actions with manually triggered comments (Action ID-1 and ID-2), manually triggered comments consistently showed higher addressing rates. This suggests that the massive feedback resulting from unconditional triggering might reduce developers’ willingness to respond. Interestingly, the LLM choice did not show high predictive importance (LLM_GPT-3.5 ranked 19/36 and LLM_GPT-4 ranked 21/36 in the feature list of the overall model; Is_Model_GPT4 ranked 23/33 in the feature list of the AI-specific model), indicating that simply employing a more advanced model does not decisively guarantee comment addressing. However, when comparing addressing rates between two model families (Table XVI), GPT-4 generated comments did show better addressing rates.
@@
For Comment Features, we find the comments were more likely to be addressed if they were concise (Text_Length, $\rho\!=\!-0.24$) or contained a higher proportion of code (Code_Text_Ratio, $\rho\!=\!0.89$), especially when they included multi-line code blocks(Has_Multiline_Code, $\rho\!=\!0.67$ not shown in the table). As shown in Table XVI, the addressing rate for both human and AI-generated comments increased noticeably when Code_Text_Ratio exceeded 0.5. This may be because multi-line code is often used to illustrate concrete solutions, allowing developers to follow comments with minimal effort through copy-pasting. Regarding topics, comments were addressed more often when they pertained to “Concurrency Control and Locking Robustness” (e.g., “using a ‘set-if-not-exists’ pattern”) (LDA_Topic_2, $\rho\!=\!0.50$), as well as “UI Usability and Accessibility Enhancements” (e.g., “adding aria-label for screen reader support”) (LDA_Topic_1, $\rho\!=\!0.41$).
@@
Within Modification Features, review comments targeting commits by experienced contributors were less likely to be addressed (Author_Prior_Commits, $\rho\!=\!-0.69$), suggesting that AI-generated reviews may be more helpful for newcomers. As shown in Table XVI, valid comments directed at project newcomers ($\textit{Author\_Prior\_Commits}\!\leq\!124$) achieved 16% addressing rate, compared to just 3.3% for the most experienced contributors. Notably, human reviews mainly focused on these newcomers, with 79% of comments directed at them, aligning with typical community practices. In contrast, automated actions reviewed code changes regardless of author experience, which may help mitigate potential experience blindspots in the review process. Finally, positive correlations for Repository Features demonstrated that comments were more likely to be addressed in large and active projects (Repo_File_Size, $\rho\!=\!0.70$; Repo_Issue_Count, $\rho\!=\!0.64$, ranked 13/36).
TABLE XIII: RQ3: Addressing rates by Trigger Mode across review actions (p-values from Fisher’s Exact Test).
Action Auto Trigger Manual Trigger p-value Total Addr. (%) Total Addr. (%) Total 2,886 9.0% 104 14.4% >>0.05 ID-1 1,595 6.8% 86 12.8% $\mathbf{\leq 0.05}$ ID-2 602 0.5% 18 22.2% $\mathbf{\leq 0.05}$ ID-3 534 25.7% ID-4 155 7.1%
TABLE XIV: RQ3: Addressing rates by LLM Series across review actions (p-values from Fisher’s Exact Test).
Action GPT-3.5 Series GPT-4 Series p-value Total Addr. (%) Total Addr. (%) Total 1,435 5.4% 1,555 12.6% $\mathbf{\leq 0.05}$ ID-1 1,431 5.5% 250 16.4% $\mathbf{\leq 0.05}$ ID-2 620 1.1% ID-3 4 0.0% 530 25.8% >0.05>0.05 ID-4 155 7.1%
TABLE XV: RQ3: Comment Addressing rates by binned Code_Text_Ratio (All Four Actions vs. Human).
Bins of Code_Text_Ratio Four Actions Human Total Addr. (%) Total Addr. (%) (–0.00, 0.00] 642 4.2% 320 71.2% (0.00, 0.11] 669 7.2% 61 65.6% (0.11, 0.26] 635 5.7% 95 75.8% (0.26, 0.52] 621 10.5% 107 78.5% (0.52, 1.00] 423 23.2% 306 89.9%
TABLE XVI: RQ3: Comment Addressing rates by binned Author_Prior_Commits (All Four Actions vs. Human).
Bins of Author_Prior_Commits Four Actions Human Total Addr. (%) Total Addr. (%) (–0.00, 30] 218 16.1% 558 80.1% (30, 124] 639 16.1% 140 82.1% (124, 319] 642 12.1% 131 71.8% (319, 1,013] 763 4.5% 32 81.2% (1013, 4,316] 728 3.3% 28 60.7%
Table XII presents the interpretability results for AI-generated comments. As indicated by the gray-shaded rows, features that rank highly in both the total and AI-specific models demonstrate the same influence direction ($\rho$). These consistent influence directions indicate that our main findings remain stable when restricting the analysis to AI-generated comments only. Although trigger-mode features become less prominent here, our earlier comparison across trigger modes (Table XVI) still supports higher addressing rates for manually triggered comments. Interestingly, when focusing solely on AI-generated comments, suggestions related to “Error Handling and Code Documentation” (e.g., “Instead of printing stack, use proper logging for exceptions”)(LDA_Topic_5, $\rho\!=\!0.58$) become more influential and are positively associated with addressing. Finally, the rise of Author_Prior_Commits to the top rank with a negative correlation ($\rho\!=\!-0.67$) highlights that there is still significant room for AI reviewers to evolve to become valuable for experienced developers.
@@
Answer to RQ3: Comment addressing is strongly associated with Source Features (human vs. AI and specific action type) and Comment Features (conciseness and code-richness are better), highlighting the importance of thoughtful automated review design.